OpenStudy (anonymous):

A chemist is mixing two solutions together. The chemist wants 100 gallons when complete. He is mixing a 24% solution with a 50% solution to get a 36% solution. How many gallons of each will he use?

OpenStudy (anonymous):

let x rep the no. of gallons of the 24% solution let y rep the no, of gallons of the 50% solution. Volume Equation 1) x + y = 100 Pure Stuff Equation 2) 0.24x + 0.50y = (0.36)(100) 3) 24x + 50y = (0.36)(100)(100) 2) x 100 Simplify 4) 24x + 50y = 3600 5) 24x + 24y = 2400 1) x 24 6) .........26y = 1200 4) - 5) 7) ............y = 1200/26 6) / 26 Substitute 7) in 1) 8) x + 1200/26 = 100 9) x = 100 - 1200/26 10) x = 2600/26 - 1200/26 11) x = 1400/26 He will use 600/13 gallons of the 50% solution abnd 700/13 of the 24% solution

OpenStudy (anonymous):

You're my hero! Thank you!

OpenStudy (anonymous):

lol anytime